1. In order to determine the identity of a particular lanthanide metal (M), a voltaic cell is constructed at 25°C with the an please answer the following question. all data and info needed is provided
Answer #1

The two half reactions are Mis) = M-3 (ag) + 3e - Cu +2 (aq) + 2e - - Curso - Let Entsm and to our are the standard reductionPutting the values P . 0.34 - E cert According to Nernst equation, the potential of the cell is given by t = t - 0.0591 log [2.36 = - 2.36 = - E° -0.00985(-2.66) M3t/m -to to.0262 Мf3) м 2 -2.32 Eo 3t/m The to above value matches Ce 3+ + 3e Ž Ce (s)